three consecutive numbers are to be found such that three times the smallest plus five times the middle number minus four times the largest is 37. find the numbers
Found 2 solutions by LinnW, Cromlix:
Answer by LinnW(1048)

Let x be the first number.
x + 1 is the second
x + 2 is the third
3( the smallest ) + 5( the middle ) -4( the largest ) = 37
3(x) + 5(x+1) -4(x+2) = 37
3x + 5x + 5 -4x -8 = 37
reorder
3x + 5x -4x + 5 -8 = 37
4x - 3 = 37
add 3 to each side
4x - 3 + 3 = 37 + 3
4x = 40
divide each side by 4
x = 10
So the numbers are 10,11,12
Contact me if you would like additional assistance.
I can work with you using an online whiteboard.

Answer by Cromlix(4381)

Hi there,
3 consecutive numbers:-
n, n + 1, n + 2.
3(n) + 5(n + 1) - 4(n + 2) = 37
3n + 5n + 5 - 4n - 8 = 37
Collect like terms.
3n + 5n - 4n = 37 - 5 + 8
4n = 40
n = 10.
3 consecutive numbers:-
10,11,12.
Hope this helps :-)